### v2

# Towards Harmonized Uncertainty Estimation for Large Language Models

Abstract
To facilitate robust and trustworthy deployment of large language models (LLMs), it is essential to quantify the reliability of their generations through uncertainty estimation. While recent efforts have made significant advancements by leveraging the internal logic and linguistic features of LLMs to estimate uncertainty scores, our empirical analysis highlights the pitfalls of these methods to strike a harmonized estimation between indication, balance, and calibration, which hinders their broader capability for accurate uncertainty estimation. To address this challenge, we propose CUE[Uncaptioned image](Corrector for Uncertainty Estimation): A straightforward yet effective method that employs a lightweight model trained on data aligned with the target LLM’s performance to adjust uncertainty scores. Comprehensive experiments across diverse models and tasks demonstrate its effectiveness, which achieves consistent improvements of up to $60\%$ over existing methods. Resources are available at https://github.com/O-L1RU1/Corrector4UE.
Towards Harmonized Uncertainty Estimation for Large Language Models
Rui Li1, Jing Long1, Muge Qi1, Heming Xia2, Lei Sha3, Peiyi Wang1, Zhifang Sui1thanks:   Corresponding author 1 Peking University 2 The Hong Kong Polytechnic University  3Beihang University o_l1ru1@stu.pku.edu.cn
1 Introduction
Uncertainty is the only certainty there is.
- by John Allen Paulos
Large Language Models (LLMs) have demonstrated exceptional capabilities in handling a wide range of downstream tasks (OpenAI, 2023; Touvron et al., 2023a, b; Dubey et al., 2024). They are gradually adopted as general-purpose API interfaces (e.g., ChatGPT [^1] https://chat.openai.com), providing valuable services and assistance in human life. Despite these impressive advancements, concerns persist regarding the tendency of LLMs to generate hallucinations and factual inaccuracies with confidence Zhang et al. (2023); Wachter et al. (2024), which may mislead users to overestimate the reliability of the information provided by these models. To mitigate this issue, uncertainty estimation (Loquercio et al., 2020) proposed quantifying the reliability of model outputs so as to ensure the robustness and trustworthiness of AI-driven services.
Harmonized uncertainty estimation is expected to encompass three key aspects: 1) Indication. The uncertainty score should clearly reflect the reliability of model responses, with higher scores signaling potential inaccuracies. This can be framed as a classification task, with “reliable” or “unreliable” as the classes. 2) Balance. Within classification framework, it’s critical to strike a balance between recall and precision, ensuring that challenging cases are appropriately flagged while minimizing the resources spent on false positives. 3) Calibration. The uncertainty score should align with human intuition and probabilistic expectations, to facilitate effective calibration. By striking a harmonized balance between these three aspects, uncertainty estimation provides an ideal measure of the model’s reliability, offering both usability and interpretability.
There has been growing interest in developing uncertainty estimation methods tailored for LLMs. However, with a thorough analysis across diverse uncertainty estimation methods, we found that there still remains a large performance gap between existing methods to achieve the harmonized uncertainty estimation. Specifically, methods that excel in one aspect fall short in others. For instance, SAR Duan et al. (2023), the outstanding and state-of-the-art method in the specific dataset SciQA Auer et al. (2023), achieves the best performance in indication but performs poorly in the view of calibration. Furthermore, we found that the combination of uncertainty scores obtained by existing methods provides little improvement in uncertainty estimation performance, suggesting that these methods are quite homogeneous. These findings highlight considerable room for refinement in uncertainty estimation.
In this paper, we introduce CUE[Uncaptioned image], a simple yet effective framework for adjusting uncertainty scores, which is orthogonal to existing uncertainty estimation methods. Specifically, we begin by curating dataset that is closely aligned with the target LLM’s performance within a particular domain of knowledge. This dataset is then utilized to train an auxiliary lightweight model, which serves as a Corrector to adjust the uncertainty scores. By integrating the Corrector trained on global alignment information with those uncertainty estimation methods that rely solely on the intrinsic logic and linguistic features of LLMs, we can significantly refine the uncertainty scores.
Our main contributions are thus as follows:
  • According to an empirical analysis of existing uncertainty estimation methods from both classification and calibration views, we found there is substantial room for improvement in their performance regarding classification indication, precision-recall balance, and calibration.
  • We propose CUE[Uncaptioned image], an uncertainty score correction framework that employs a classifier, aligned with the model’s task performance, as a Corrector to adjust uncertainty scores. This Corrector allows for seamless integration with existing uncertainty estimation methods.
  • Extensive experiments and explorations are also conducted in areas such as generalization. The results demonstrate that our CUE[Uncaptioned image] consistently enhances various existing uncertainty estimation methods, showing significant improvements in a harmonized manner across diverse data domains and target models.
2 Related Work
2.1 Uncertainty Estimation for LLMs
As illustrated in Figure 1, uncertainty estimation methods for LLMs can be broadly categorized into logit-based methods, verbalized methods, consistency-based methods and internal state-based methods.
Logit-based methods are the most widely used and effective approaches in uncertainty estimation. Predictive Entropy (PE) (Malinin and Gales, 2020) defined uncertainty as the entropy of the output logits distribution, which is widely adopted and built upon in subsequent research. Following that, Kuhn et al. (2023) introduced semantic entropy (SE) that estimates uncertainty by marginalizing over semantically-equivalent samples in NLG tasks. Duan et al. (2023) proposed Shifting Attention to Relevance (SAR), which focuses on relevant information and assigns significance weights to tokens based on their contributions to the overall response. Yaldiz et al. (2024) introduced a Learnable Response Scoring Function (LARS), which utilizes supervised data to capture complex token-probability dependencies.
Figure 1: A concise overview figure of various uncertainty estimation method categories, including logit-based methods, verbalized methods, consistency-based methods, and internal state-based methods.
Figure 2: The performance of existing uncertainty estimation methods, evaluated on the SciQA dataset with the LLaMA-3-8B-Instruct model as the target, and the improvements after applying the Corrector. Note that a lower ECE score indicates better performance, so we report its reduction.
Figure 3: AUROC improvement across uncertainty scores combination from different existing methods.
Verbalized methods (Xiong et al., 2023; Groot and Valdenegro-Toro, 2024) leverage LLMs’ strong language and instruction-following abilities to express uncertainty, often by prompting the model to provide an uncertainty score. However, studies (Ni et al., 2024; Madhusudhan et al., 2024; Becker and Soatto, 2024) have shown that LLMs struggle with faithfully conveying their uncertainties, particularly due to overconfidence. Consistency-based methods, such as those proposed by Li et al. (2024b) and Becker and Soatto (2024), assess uncertainty through multiple generated answers, using techniques like perturbation and aggregation to improve reliability. Pedapati et al. (2024) further reduced overconfidence by guiding LLMs to justify their answers. Internal state-based methods (Azaria and Mitchell, 2023; Liu et al., 2024) analyze LLM activations to predict errors, with Kadavath et al. (2022) and Ji et al. (2024) exploring self-evaluation and probing estimators to enhance uncertainty estimation.
Due to space limitations, a more detailed discussion of related work is provided in the Appendix LABEL:app:more_work.
3 Preliminary Study
Figure 4: An overview of uncertainty score correction framework. Firstly, we construct a dataset that closely aligns with the target model’s performance. This dataset is then utilized to train a lightweight auxiliary model that serves as a correction module, enabling seamless integration with existing uncertainty estimation methods to produce corrected uncertainty scores.
3.1 Limitation of Existing UE Methods
We evaluate existing UE methods from both classification and calibration views, focusing on three key aspects of uncertainty scores: indication, balance, and calibration. From the classification view, uncertainty scores are utilized to guide the classification process. Instances with scores above a threshold are classified as $c_{1}$ (unreliable) and those below as $c_{0}$ (reliable). We employ AUROC to measure how well the scores indicate unreliability and F1 score to evaluate their balance between precision and recall. The calibration view involves a more rigorous assessment and interpretation of uncertainty scores. Well-calibrated scores should align with human probabilistic intuition and provide more precise instance rankings. We use ECE to assess calibration.
Basic methods exhibit poor indication performance.
Firstly, we focus on representative but naive methods including Lexical Similarity (LS) (Fomicheva et al., 2020), Verbal Confidence (VC) (Xiong et al., 2023), P(true) (Kadavath et al., 2022), and Predictive Entropy (PE) (Malinin and Gales, 2020) that belong to four categories: consistency-based methods, verbal confidence methods, internal state-based methods, and logit-based methods, respectively. As shown in Figure 2 and Table 1, the AUROC scores for these methods across the target models and datasets exhibit general low performance, which is even close to random guessing.
Enhanced logit-based methods typically have low F1 scores.
Some enhanced methods such as Length-normalized Predictive Entropy (LN-PE) (Malinin and Gales, 2020), SAR-t, SAR-s, SAR, and Semantic Entropy (SE) (Kuhn et al., 2023), make tailored adjustments to refine predictive entropy process, which show improvements over PE in terms of AUROC. However, no one is universally optimal for all target models and datasets. Moreover, as depicted in Figure 2 and Figure 5, the F1 scores of those methods are particularly low. This indicates that although those methods provided uncertainty scores with some potential to indicate the reliability of model response, they still fall short in striking a balance between precision and recall.
Most existing methods fall short in calibration.
As shown in Table 1 and Figure 6, it appears that prior methods have overlooked the calibration aspect, resulting in relatively poor performance in terms of ECE scores.
3.2 Inter-method Cooperation
We examined whether the uncertainty scores derived from one uncertainty estimation method could refine the scores obtained from another method. Specifically, we integrated the uncertainty scores from each method using the weighted combination and compared its performance with the top-performing method in the pair. As illustrated in Figure 3, these integrations do not enhance overall performance and may lead to a decline. This underscores the limitations in the complementary nature of existing methods.
Our analysis reveals a significant performance gap among existing methods in achieving harmonized uncertainty estimation, as individual methods excel in specific aspects but underperform in others. Furthermore, combining uncertainty scores from different methods yields minimal to no improvement, underscoring their homogeneous and non-complementary nature.
4 Method
In this section, we introduce CUE[Uncaptioned image], a correction framework featuring an intuitive approach to directly optimizing for uncertainty estimation, where a Corrector is trained using a lightweight model to refine the uncertainty score. Through this method, we provide a more robust solution for uncertainty estimation. As shown in Figure 4, Our method comprises three main steps including dataset crafting, corrector training and uncertainty correcting.
4.1 Dataset Crafting
[…]
To mitigate false positives, we apply rigorous thresholds and strict prompting rules. The final judgment is determined using an “OR” logic: $\mathcal{M}(r_{i},a_{i})=\mathcal{M}_{\text{Rule}}(r_{i},a_{i})\lor\mathcal{M}_{\text{LLM}}(r_{i},a_{i})$, preventing the omission of positive instances.
After that, a binary label $c_{i}$ is assigned to each sample, defined as
$c_{i}\text{ = }\mathcal{M}(r_{i},a_{i})$ (1)
By pairing question $q_{i}$ with the label $c_{i}$, we form a correction dataset $\mathcal{D}_{\text{cor}}=\{(q_{i},c_{i})\mid i=1,\ldots,n\}$, which serves as a representation of the target model’s performance in generating correct responses across a particular knowledge domain. To directly associate the questions with uncertainty, we transform the dataset form into $\mathcal{D^{*}}_{\text{cor}}=\{(q_{i},1-c_{i})\mid i=1,\ldots,n\}$.
4.2 Corrector Training
Employing the correction dataset $\mathcal{D^{*}}_{\text{cor}}$, we train a classifier to align with the performance of the target model. Specifically, the classifier integrates a fully connected layer following a lightweight encoder model, such as RoBERTa (Liu, 2019) and Deberta (He et al., 2021a), with the representation of the special token $[CLS]$ as its input, denote as $\mathbf{h}_{[CLS]}\in\mathbb{R}^{d}$. The output of the classifier is given by $\hat{y}_{i}=\sigma\left(\mathbf{W}\cdot\mathbf{h}_{[CLS]}+b\right),$ where $\sigma(z)$ is the sigmoid function, used to compute the likelihood $y_{i}$ that a data point belongs to label $c_{1}$. During training, we minimize the binary cross-entropy loss function $\mathcal{L}=-\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ across the correction dataset.
This results in a Corrector, an auxiliary component that can be integrated with existing uncertainty estimation methods to enhance their reliability.
4.3 Uncertainty Correcting
We derive the probability that an instance $x$ belongs to category $c_{1}$ from the Corrector. This probability, denoted as the correction score $C(x)$, can be utilized to adjust the uncertainty scores to align with the target model’s performance, thereby refining the uncertainty estimation process.
In the refinement process, we first normalize the uncertainty scores generated from existing UE methods to match human probabilistic intuition, ensuring they fall within the range $[0,1]$. Normalization is achieved via $U_{\text{norm}}(x)=\frac{U(x)-\min(U)}{\max(U)-\min(U)}$, where $U(x)$ represents the uncertainty score for a specific instance $x$, computed by a chosen UE method. The terms $\min(U)$ and $\max(U)$ denote the minimum and maximum uncertainty scores across the entire dataset, respectively. Following normalization, we apply our correcting by combining the normalized score $U_{\text{norm}}(x)$ with the correction score $C(x)$ generated by the Corrector. The combination employs a weighted approach, where the corrected uncertainty score $U_{\text{cor}}(x)$ is computed as:
$U_{\text{cor}}(x)=w^{*}\cdot U_{\text{norm}}(x)+(1-w^{*})\cdot C(x)$ (2)
The optimal weight $w^{*}$ is determined through a grid search on the development dataset. This weighted method ensures that the corrected uncertainty scores balance the contributions of both the original and correction scores, thereby enhancing the reliability of the uncertainty estimation.
5 Experiments
5.1 Experiments Setup
5.1.1 Models
Target models
We selected the OPT-6.7B222huggingface.co/facebook/opt-6.7b. (Zhang et al., 2022), a model widely utilized in previous studies (Kuhn et al., 2023; Duan et al., 2023), and the advanced open-source model LLaMA-3-8B-Instruct333huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. (Dubey et al., 2024) as the target models for our main experiments.
Base Models
We employed lightweight encoder models as the base model to train the Corrector, including models from the RoBERTa series Liu (2019) and DeBERTa series He et al. (2021a, b).
5.1.2 Metrics
AUROC
We use the area under the receiver operating characteristic curve (AUROC) to evaluate uncertainty estimation methods from a classification view. In our setting, an AUROC of 1 signifies perfect indicative performance to distinguish between samples the target model can answer reliably and those it cannot, while an AUROC of 0.5 indicates that the estimation is no better than random guessing.
F1 Score
F1 score is used to evaluate the balance between precision and recall in classification tasks. It is the harmonic mean of precision and recall, where both are equally important. The F1 score ranges from 0 to 1, with 1 indicating perfect precision and recall.
[…]
ECE
We use Expected Calibration Error (ECE) to evaluate the performance of calibration, which is calculated by partitioning predicted confidence scores into bins and comparing the average confidence in each bin to the actual fraction of correct predictions, formalized as
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (4)
In the computing of ECE, we treat the confidence score as 1 minus uncertainty score.
5.1.3 Datasets
We focus on the question-answering task using two representative datasets in the main experiments: TriviaQA (Joshi et al., 2017), and SciQA (Auer et al., 2023). TriviaQA comprises 95,000 question-answer pairs created by trivia enthusiasts, supplemented with independently sourced evidence documents. SciQA contains 2,565 question-answer pairs fetched from the open research knowledge graph, covering several research fields ranging from science and technology like Computer Science, Engineering, Chemistry, and Geology, life sciences like Immunology and Genetics to social sciences like Economics and Urban Studies.
5.1.4 Baselines
We select a variety of representative uncertainty estimation methods as baselines, with a particular focus on logits-based methods.
Among the baselines, we cover multiple categories, including logit-based, verbalized, internal state-based, and consistency-based methods, including: Lexical Similarity (LS) (Fomicheva et al., 2020), which computes the similarity between multiple sentences as a measure of consistency; Verbal Confidence (VC) (Xiong et al., 2023), which requires the target model to respond and provide a confidence score; P(True) (Kadavath et al., 2022), which first asks the target model to propose an answer and then evaluates it using an internal probability mechanism; and Predictive Entropy (PE) (Malinin and Gales, 2020), which calculates uncertainty by measuring the entropy of the predictive posterior.
We also explore a series of advanced logit-based methods including: Length-normalized Predictive Entropy (LN-PE) (Malinin and Gales, 2020), which adjusts PE by normalizing it according to sentence length; Semantic Entropy (SE) (Kuhn et al., 2023), which clusters sentences with equivalent meanings and calculating cluster-wise entropy; and Shifting Attention to Relevance (SAR) (Duan et al., 2023), which encompasses SAR-t, SAR-s and SAR, donated as the token-shifted predictive entropy, sentence-shifted predictive entropy, and both token- and sentence-shifted predictive entropy respectively.
TriviaQA SciQA
AUROC($\uparrow$) ECE($\downarrow$) AUROC($\uparrow$) ECE($\downarrow$)
Method Vanilla +Corrector Improv Vanilla +Corrector Improv Vanilla +Corrector Improv Vanilla +Corrector Improv
OPT-6.7B
LS 46.49 65.11 +18.62 72.71 41.76 -30.94 44.12 49.40 +5.29 76.38 32.78 -43.60
VC 60.41 70.55 +10.15 49.13 27.61 -21.52 51.69 56.55 +4.86 62.65 38.99 -23.66
P(True) 66.74 72.29 +5.84 45.00 32.63 -12.80 56.12 59.49 +3.37 58.79 34.52 -24.27
PE 56.36 66.62 +10.25 42.39 20.28 -22.12 50.07 56.02 +5.95 62.05 36.92 -25.13
LN-PE 78.37 79.93 +1.57 32.29 20.80 -11.49 60.88 64.23 +3.35 49.52 34.68 -14.84
SE 80.66 81.00 +0.34 36.64 27.05 -9.59 64.52 66.15 +1.63 52.66 42.23 -10.43
SAR-t 78.24 80.21 +1.97 40.14 37.85 -2.30 60.00 63.75 +3.74 45.33 44.19 -1.14
SAR-s 51.77 55.83 +4.06 53.78 49.65 -4.13 53.20 54.15 +0.95 76.21 34.83 -41.38
SAR 75.32 78.67 +3.35 40.61 31.02 -9.59 60.04 62.72 +2.68 49.40 38.99 -10.41
LLaMA-3-8B-Instruct
LS 19.57 69.82 +50.25 70.25 7.41 -62.84 53.67 65.38 +11.71 38.64 18.19 -20.45
VC 62.34 74.89 +12.55 23.41 16.78 -6.63 68.22 72.15 +3.93 31.88 19.47 -12.36
P(True) 57.14 72.29 +15.15 24.67 19.84 -4.83 65.63 71.41 +5.78 34.56 31.92 -2.64
PE 64.52 69.76 +5.25 21.38 17.24 -4.13 66.54 67.98 +1.44 40.67 34.07 -6.60
LN-PE 72.55 74.79 +2.24 14.31 11.53 -2.79 69.48 71.56 +2.08 29.38 23.76 -5.62
SE 80.92 82.12 +1.20 13.07 12.76 -0.31 71.59 72.93 +1.34 30.54 25.23 -5.30
SAR-t 79.55 79.93 +0.38 16.40 13.70 -2.70 72.26 73.87 +1.61 30.37 26.81 -3.56
SAR-s 69.87 77.09 +2.95 23.17 20.00 -3.17 74.96 75.72 +0.76 38.54 36.18 -2.37
SAR 80.92 81.90 +0.98 16.17 13.76 -2.41 73.88 75.19 +1.31 28.97 25.60 -3.37
Table 1: AUROC and ECE scores (%) on the TriviaQA and SciQA datasets obtained by applying the Corrector to existing uncertainty estimation methods. LS denotes the Lexical Similarity method. VC denotes the Verbal Confidence method. PE denote the Predictive Entropy method. LN-PE denotes the Length-normalized Predictive Entropy method. SE denote the Semantic Entropy. SAR-t refers to the token-level version of the SAR method, while SAR-s denotes the sentence-level version.
5.1.5 Implementation Details
Dataset Splitting
For the TriviaQA dataset, we randomly selected 5,000 samples from the training set for data crafting and corrector training. For datasets with limited data, SciQA, we utilized the entire training set. We then used half of the test set to search for the optimal hyperparameter w, while the other half was employed to evaluate the method’s effectiveness.
Hyperparameter
For each dataset and model pair, we train a corresponding Corrector, which is universally applicable across various methods. Additionally, for every method, dataset, and model combination, we derive the weight using the development set respectively. The sensitivity analysis of hyperparameter $w^{*}$ and its configuration within the cross-domain experiments are elaborated upon in Appendix LABEL:app:detail_of_w.
Figure 5: The performance gains of using the Corrector to adjust the uncertainty scores for various methods on the datasets of TriviaQA and SciQA, and the target models of LLaMA-3-8B-Instruct and OPT-6.7B, are evaluated in terms of F1 score.
5.2 Main Result
We has evaluated existing methods in Section 4 and found that there still remains a large performance gap between existing methods to achieve the harmonized uncertainty estimation. In this part, we present the performance of CUE[Uncaptioned image] from both classification and calibration views, demonstrating that integrating a Corrector with existing UE methods significantly enhances uncertainty estimation across multiple dimensions, including classification indication, precision-recall balance, and calibration.
Classification View
As illustrated in Table 1, the Corrector has resulted in significant improvements, with an average AUROC score increase of 0.27 for TriviaQA and 0.09 for SciQA. Even when applied to challenging methods such as SE and SAR, the Corrector boosts AUROC scores by 0.01 to 0.03. Since AUROC reflects the UE method’s ability to assign higher scores to instances where the target model responds unreliably compared to those it responds to reliably, these improvements indicate that the deployment of the Corrector enhances the overall indicative capacity of the uncertainty scores, making it more effective for users in determining whether to trust the model’s responses.
Furthermore, as illustrated in Figure 5, the F1 score is also boosted by the Corrector, achieving an average increase of 38.97%. This notable improvement demonstrates the Corrector’s ability to help balance precision and recall, effectively mitigating the polarization tendency in the uncertainty scores observed in previous methods.
Calibration View
Although calibration is not the direct training objective of our Corrector, its application yields favorable calibration results. When employing the OPT-6.7B model as the target, we observed average ECE reductions of 0.34 on TriviaQA and 0.21 on SciQA. With the LLaMA-3-8B-Instruct model as the target, the reductions are 0.11 and 0.07, respectively—still considerable. To further illustrate the calibration performance facilitated by the Corrector, we provide calibration plots in Figure 6.
We also conducted extensive experiments on the generalization performance of the Corrector. Due to space constraints, we put the detailed results and analysis in the Appendix LABEL:app:gene.
In summary, integrating the Corrector helps achieve harmonized uncertainty estimation. With the Corrector, we can improve the reliability of uncertainty scores and alignment with the actual performance of the model. The analysis of the pure Corrector’s performance can be found in Appendix LABEL:app:pure_corrector.
Figure 6: Calibration Plots. These plots depict the relationship between predicted confidence and observed frequencies. The diagonal line represents perfect calibration, where predicted confidence aligns precisely with actual outcomes. Bars extending above the diagonal indicate underestimation of confidence, while bars below the diagonal reflect overestimation. The final plot highlights the optimal calibration performance achieved through our Corrector.
5.3 Ablation Study
We conducted ablation studies to scrutinize the impact of the base model, the correction score formats and its acquisition methods.
Formats
We compared the efficacy of probabilistic values versus label values for correction. As shown in Table 2, probabilistic correction scores demonstrate clear superiority, as they allow finer-grained adjustments by leveraging a broader spectrum for integration. Conversely, discrete values, such as 0 and 1, tend to introduce significant biases in the corrected uncertainty scores.
Base Model
We utilized various encoder models as base models to train the Corrector and assess the impact on correction performance. Specifically, we employed models from the RoBERTa series, including RoBERTa-base444huggingface.co/FacebookAI/roberta-base and RoBERTa-large555huggingface.co/FacebookAI/roberta-large, as well as models from the DeBERTa series, including DeBERTa-base666huggingface.co/microsoft/deberta-base, DeBERTa-v3-large777huggingface.co/microsoft/deberta-v3-large, DeBERTa-v3-base888huggingface.co/microsoft/deberta-v3-base, DeBERTa-v3-small999huggingface.co/microsoft/deberta-v3-small, and DeBERTa-v3-xsmall101010huggingface.co/microsoft/deberta-v3-xsmall. These models represent different types, series, and sizes. As illustrated in Figure 7, more advanced, later-generation, and larger models yield superior results.
Figure 7: The overall AUROC score gains achieved by Correctors trained on different base models across various UE methods on the SciQA dataset and Llama-3-8B-Instruct target model.
Acquisition
We compare correction scores from a lightweight classifier with those estimated using GPT-4o. We attempt not to rigorously assess the target model’s answers but to predict its reliability. Despite GPT-4’s strong performance in question answering, our results show it is less effective than the classifier in directly predicting reliability of target models when faced with questions. Additionally, as detailed in Section 3.2, combining uncertainty scores from different UE methods does not improve and may even degrade performance. This highlights the Corrector’s unique role as a complement to existing UE methods.
We also conducted experiments to compare our Corrector with other supervised methods specifically designed for uncertainty estimation. The experimental results demonstrated that our supervised pipeline offered significant advantages in enhancing uncertainty estimation. Detailed results and analysis can be found in Appendix LABEL:app:comparative.
Methods AUROC ($\uparrow$) ECE ($\downarrow$)
Corrector 69.87 6.73
Original Best 80.92 11.53
+Corrector Probability 82.12 10.46
+Corrector Label 80.92 11.53
+GPT-4o Score 80.92 11.53
Table 2: Ablation Study. LLaMA-3-8B-Instruct as the target model and TriviaQA as the test dataset. Original Best refers to the peak performance achieved by various baseline when the Corrector is not incorporated.
6 Conclusion
Our study highlights the limitations of current uncertainty estimation methods in terms of classification accuracy, precision-recall balance, and calibration. We introduce an innovative uncertainty score correction framework that utilizes a classifier as a Corrector to refine these scores, ensuring alignment with the model’s true task performance. This Corrector integrates seamlessly with existing methods, enhancing their effectiveness. Extensive experiments validate that the Corrector consistently improves performance across various metrics, data domains, and target models. Furthermore, our ablation study underscores the Corrector’s capacity to provide substantial and heterogeneous improvements to existing uncertainty estimation techniques.
Limitations
Although the CUE method proposed in this paper demonstrates good performance, its dependence on labeled data and its generalization ability across different data domains and target models may be limitations. We only compared our method with works that have open-source code, which are often designed for white-box models. Therefore, the effectiveness of our method on black-box models has not been demonstrated through experiments. However, our method does not necessitate access to the inner states of target models, making it a general enhancement strategy for both black-box and white-box uncertainty estimation.
Ethics Statement
In this study, we introduce a method for improving uncertainty estimation in the context of LLMs, which presents no immediate ethical concerns, but certain considerations must be addressed. Uncertainty estimation has significant potential to evaluate the reliability and safety of LLM outputs. However, this potential benefit comes with the risk that systematic mistakes in the uncertainty assessment could foster unfounded and misplaced confidence. Consequently, even re-calibrated uncertainty estimates should be interpreted cautiously, particularly in critical decision-making scenarios where the consequences of inaccuracies can be profound.
The datasets used in our experiment are publicly released and labeled through interaction with humans in English. In this process, user privacy is protected, and no personal information is contained in the dataset. The scientific artifacts that we used are available for research with permissive licenses. And the use of these artifacts in this paper is consistent with their intended use. Therefore, we believe that our research work meets the ethics of ACL.
Acknowledgements
This paper is supported by NSFC project 62476009.
Appendix A Appendix